### v2

# Advanced Video Inpainting Using Optical Flow-Guided Efficient Diffusion

Abstract
Recently, diffusion-based methods have achieved great improvements in the video inpainting task. However, these methods still face many challenges, such as maintaining temporal consistency and the time-consuming issue. This paper proposes an advanced video inpainting framework using optical Flow-guided Efficient Diffusion, called FloED. Specifically, FloED employs a dual-branch architecture, where a flow branch first restores corrupted flow and a multi-scale flow adapter provides motion guidance to the main inpainting branch. Additionally, a training-free latent interpolation method is proposed to accelerate the multi-step denoising process using flow warping. Further introducing a flow attention cache mechanism, FLoED efficiently reduces the computational cost brought by incorporating optical flow. Comprehensive experiments in both background restoration and object removal tasks demonstrate that FloED outperforms state-of-the-art methods from the perspective of both performance and efficiency.
footnotetext: *Work during DAMO Academy internship. $\dagger$ Co-corresponding author.
1 Introduction
Video inpainting aims to predict corrupted regions by filling them with contextually appropriate and temporally coherent content, which plays an essential role in computer vision, particularly in background restoration (BR) [23, 17, 34] and object removal (OR) [4, 33, 19].
Although conventional transformer-based approaches [15, 5, 37] have made great progress on standard benchmarks [31, 20], they encounter difficulties when requiring the synthesis of new content not present in existing frames. Currently, diffusion models [36, 38] have demonstrated significant prowess in text-guided video inpainting. However, temporal consistency remains an area with substantial scope for further improvement. And current methods struggle to generate satisfying results on BR and OR tasks, which is probably caused by the training priors of powerful base models. As demonstrated in Fig. 1, contents inpainted by CoCoCo [38] lacks harmony with its surroundings. This disharmony can be alleviated by providing motion guidance that is aligned with the surroundings. Optical flow, a key modality for motion information, provides significant guidance and enhances temporal consistency. Notably, integrating optical flow into video inpainting requires additional operations, including flow completion and effective incorporation, which incur extra computational costs. And diffusion inherently suffers from efficiency due to the multi-step denoising process. Consequently, when leveraging optical flow, it is essential to consider efficiency enhancements tailored to the multi-step characteristics of diffusion models.
Figure 1: Superiority of FloED on background restoration and object removal.
Based on the above analysis, we propose an advanced video inpainting framework using optical Flow-guided Efficient Diffusion, called FloED. By leveraging motion information, our approach aims to enhance the reliability and performance of diffusion-based techniques in BR and OR applications. Initially, built on the large-scale T2I model, we inflate the origin layers and integrate pre-trained motion module [7] to establish a basic text-to-video framework. Then we fine-tune the motion module to effectively adapt it for the video inpainting domain. By adopting this framework as the primary inpainting branch, our innovations focus on three key aspects: (1) We design a tailored flow branch that completes corrupted flow while maintaining consistent channel numbers with the primary branch. Next, we integrate a multi-scale flow adapter that incorporates flow features into the decoder blocks of the primary U-Net architecture, enabling FloED to utilize motion information more effectively. (2) Building on the observation that adjacent latents share similar motion patterns [35] and recognizing that the essence of diffusion models involves fundamentally multi-step sampling processes, we introduce a training-free latent interpolation technique. This method leverages warping operation guided by optical flow to effectively accelerate the multi-step denoising process during early denoising stage. Furthermore, by incorporating a flow attention cache mechanism during the remain denoising stage as a complementary speed-up solution, we efficiently minimize the additional computational burden typically introduced by flow adapters and flow branches. (3) Recognizing that state-of-the-art image inpainting models significantly outperform video inpainting diffusion models, we utilize an anchor frame strategy to enhance the quality of video inpainting outcomes.
Currently, there is no comprehensive benchmark for evaluating diffusion-based generative approaches in video inpainting. This deficiency presents a substantial challenge, as it limits the ability to rigorously assess and compare the efficacy of various inpainting methodologies. To bridge this gap, we have developed an extensive benchmark that meticulously encompasses both BR and OR tasks. Our main contributions are as follows:
  • We propose a dedicated dual-branch architecture that incorporates motion guidance with a multi-scale flow adapter, thereby enhancing temporal consistency and overall quality.
  • We introduce a training-free latent interpolation technique that leverages optical flow to speed up the multi-step denoising process. Complemented by a flow attention cache mechanism, FloED efficiently reduces the additional computational costs introduced by the flow.
  • We conducted extensive experiments, including both quantitative and qualitative evaluations, to validate that FloED demonstrates superiority over other state-of-the-art methods in both BR and OR tasks.
2 Related Work
2.1 Conventional Video Inpainting
Conventional video inpainting mainly focuses on two types of tasks: object removal and background restoration. Object removal [4, 33, 19] aims to eliminate unwanted objects from video frames. And background restoration [23, 17, 34] involves seamlessly reconstructing missing region of the background with coherent content.
Given the fact that it’s more simple to complete flows instead of directly filling masked regions [32], conventional transformer-based models [15, 35, 37] primarily rely on optical flow to propagate features or pixels for better inpainting outcomes. Specifically, E2FGVI [15] introduces an end-to-end framework with flow-guided feature propagation. FGT [35] combines decoupled spatiotemporal attention with a flow-guided content propagation. ProPainter [37] advances this field by merging dual-domain propagation with a mask-guided transformer.
Although propainter [37] has shown significant improvements on standard benchmarks [20, 31], it still struggles with generating coherent new content that isn’t present in existing frames. There is a pressing need to develop more robust solutions.
2.2 Diffusion-based Video Inpainting
In recent years, diffusion models [22, 9] have revolutionized the field of content generation, showcasing exceptional abilities in creating highly realistic outputs [21, 2]. The development of text-to-video (T2V) [7, 10] and image-to-video (I2V) [30, 3] diffusion generative models, has paved the way for innovations in text-guided video inpainting. Models like AVID [36] and CoCoCo [38], have highlighted the routine practice of employing pre-trained Stable Diffusion Inpainting models [2] equipped with motion modules [7] for text-guided video inpainting. Specifically, AVID [36] introduce zero-shot generation pipeline with a structure guidance module. CoCoCo [38] implements enhanced attention mechanisms to address text-video alignment and ensure motion consistency across frames. While these advancements have contributed to progress, existing diffusion-based video inpainting methods still exhibit significant room for improving temporal consistency and they perform suboptimally on both OR and BR tasks. Compared with them, FloED showcases distinct advantages over temporal consistency and overall quality.
Figure 2: Overview of FloED. FloED employs a dual-branch architecture implemented through a two-stage training approach. In the first training stage, we focus exclusively on the upper branch, optimizing the motion layer to adapt specifically to the video inpainting domain. Subsequently, we introduce a dedicated flow branch complemented by a multi-scale flow adapter, which provides flow guidance covering upblocks of primary UNet. During the inference phase, we enhance efficiency by integrating the flow attention cache (right part).
3 Preliminaries
LDM [21] leverages a pre-trained Variational Autoencoder (VAE) to operate in the latent space instead of pixel space. The diffusion forward process is imposing noise on a clean latent $\mathbf{z}_{0}$ for $T$ times. A property of the forward process is that it admits sampling $\mathbf{z}^{t}$ at random timestep $t$:
$q(\mathbf{z}^{t}|\mathbf{z}^{0})=\mathcal{Q}(\mathbf{z}^{0},t)=\mathcal{N}(\mathbf{z}^{t};\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (1)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$, $\beta_{s}$ is the variance schedule for the timestep $s$. The backward process applies a trained UNet $\epsilon_{\theta}$ for denoising: $p_{\theta}(\boldsymbol{z}_{t-1}|\boldsymbol{z}_{t})=\mathcal{N}(\boldsymbol{z}_{t-1};\mu_{\theta}(\boldsymbol{z}_{t},t),\Sigma_{\theta}(\boldsymbol{z}_{t},t)),$ where distribution parameters $\mu_{\theta}$ and $\Sigma_{\theta}$ are computed by the denoising model $\theta$. To train a conditional LDM, the objective is given by:
[…]
where $\epsilon_{\theta}(\mathbf{z}_{t},t,c)$ is the predicted noise based on $\mathbf{z}_{t}$, the time step $t$ and the condition $c$. Once trained, we could leverage the deterministic sampling of DDIM [9] and PNDM [16] to denoise $\mathbf{z}_{t}$:
$\mathbf{z}_{t-1}$ $=\sqrt{\alpha_{t-1}}\underbrace{\hat{\mathbf{z}}_{t\rightarrow 0}}_{\text{predicted `$\mathbf{z}_{0}$'}}+$ (3)
$\underbrace{\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(\mathbf{z}_{t},t,c)}_{\text{direction pointing to $\mathbf{z}_{t}$}}+\underbrace{\sigma_{t}\epsilon_{t}}_{\text{random noise}},$
where $\sigma_{t}$ are hyper-parameters. The term ${\mathbf{z}}^{t}_{t\rightarrow 0}$ represents the predicted $\mathbf{z}_{0}$ at time step $t$, For conciseness and to circumvent any potential confusion with the concept of optical flow, we subsequently refer to $\hat{\mathbf{z}}_{t\rightarrow 0}$ as $\hat{\mathbf{z}}_{0}$. The precise formulation is as follows:
$\hat{\mathbf{z}}_{t\rightarrow 0}\ =\mathcal{P}(\mathbf{z}_{t},\epsilon_{\theta})=(\mathbf{z}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{z}_{t},t,c))/\sqrt{\alpha_{t}}.$ (4)
4 Methods
Building upon Stable Diffusion inpainting model [1], we integrate the motion module from Animatediff [7], adopting it as primary inpainting branch of our framework. As illustrated in Fig. 2, the training process of FloED can be divided into two stages. Initially, we adapt primary branch to video inpainting by fine-tuning the motion modules (Sec. 4.1). Our approach then incorporates a dedicated flow branch to complete corrupted flows estimated from masked frames, alongside a multi-scale flow adapter that provides motion guidance for the primary inpainting branch (Sec. 4.2). To further enhance video inpainting results, we implement an anchor frame strategy to leverage the priority of the image inpainting diffusion model [2] (Sec. 4.3). Furthermore, we introduce a training-free denoising acceleration technique that leverages optical flow for latent interpolation, compensated with a flow attention caching mechanism in the reference phase. We substantially enhance efficiency while significantly reducing the additional computational overhead the flow introduces. (Sec. 4.4).
4.1 Domain Adaptation
Given an original video sequence $\mathbf{V_{0}}=\{\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{N-1}\}\in\mathbb{R}^{N\times 3\times H\times W}$ and a binary mask sequence $\mathbf{m}=\{\mathbf{m}_{0},\mathbf{m}_{1},\ldots,\mathbf{m}_{N-1}\}\in\mathbb{R}^{N\times 1\times H\times W}$, corrupted frames $\mathbf{V_{m}}$ are obtained by applying the Hadamard product as follows: $\mathbf{V_{m}}=\mathbf{V_{0}}\odot\mathbf{m}$. We aim to generate a set of spatiotemporally consistent inpainted outcomes with temporally coherent and contextually appropriate content in the corrupted area while ensuring out-of-mask areas are unchanged.
Following previous solutions [36, 38], FloED is built on Stable Diffusion (SD) inpainting model [1], equipped with motion module from AnimateDiff-v3 [7] for better temporal consistency. Notably, during the first stage, only the primary branch is taken into consideration, and we only optimize the motion layer for adapting to the video inpainting.
Figure 3: Illustration of flow-guided latent interpolation (left) and warping operation (right) during the denosing process.
4.2 Flow Guidance in Video Inpainting
Typically, for video inpainting methods [15, 5, 37] that utilize optical flow, it is typically necessary to first complete the corrupted flow because only an estimated flow from the masked frames is initially available. Specifically, we first employ a pre-trained flow estimator [24] to perform corrupted flow estimation from $\mathbf{V_{m}}$. Subsequently, we introduce a flow sub-branch composed of the first layer from each block of the primary branch, which is independent of the timestep. As illustrated in Fig. 2, the flow sub-branch completes the corrupted flow while maintaining consistent channel numbers with the latent representations of the corresponding frames in the primary branch. This approach enables more effective utilization of motion information by integrating multi-scale flow adapters into all upblocks of the primary UNet. As for the flow adapter, it consists of learnable modules with decoupled cross-attention mechanisms that embed flow features into the primary branch (Fig. 2), thereby providing robust motion guidance effectively. We choose to insert the flow adapter between the text attention and the motion layer because it can compensate for the information provided by the prompt and further supply temporal guidance to the motion layer. Consequently, in the second training phase, we also trained the motion layer simultaneously. An additional optical loss $\mathcal{L}_{\textrm{flow}}$ is introduced in the second stage: ($F$ represents the ground truth flow.)
$\mathcal{L}_{\textrm{sec}}=\mathcal{L}_{\textrm{diff}}+\lambda*\mathcal{L}_{\textrm{flow}},\mathcal{L}_{flow}=\|\hat{F}-F\|_{1}$ (5)
4.3 Anchor Frame Strategy
Recognizing that state-of-the-art image inpainting models significantly outperform their video inpainting counterparts, we introduce an anchor frame strategy to enhance the quality of video inpainting results. As illustrated in Fig. 2, for a given video sequence $\mathbf{V_{0}}$, we select an additional frame from the beginning of the sequence to serve as an anchor frame. We then utilize a pre-trained text-to-image (T2I) inpainting model [2] to reconstruct its corrupted region in advance. Subsequently, we concatenate the inpainted anchor frame with the noised video frames $\mathbf{V_{t}}$. This approach provides additional texture guidance to the video frames during the denoising process. After denoising, the anchor frame is discarded. This strategy leverages the superior performance of image inpainting models to improve the overall quality of video inpainting.
Figure 4: Qualitative Results. The orange dotted line indicates the input source.
4.4 Efficient Inference for FloED
Based on multi-step sampling processes of diffusion, we further propose a training-free latent interpolation technique that leverages optical flow to speed up the denoising process. This approach is complemented by a flow attention cache mechanism during the inference phase.
Optical Flow Attention Cache Unlike the primary branch, the flow sub-branch is independent of timestep. During the inference phase, we utilize the flow sub-branch exclusively for flow completion in the first step and then use these completed flows for all subsequent steps. Regarding the flow adapter, it introduces additional computations by calculating the flow attention at every denoising step and multiple resolutions. To optimize this process, we establish a cache mechanism by computing the keys and values only during the first step and storing them in the memory bank (right part of Fig. 2). For the remaining steps, the cached keys and values are directly retrieved from the memory bank, eliminating the need for repeated calculations and enhancing efficiency.
Training-Free Denoising Speed-up Since adjacent feature latent exhibit similar motion patterns [35] and diffusion models generate high-level content early in the denoising process [13], we aim to speed-up the denoising process by interpolating latent using the completed flow. Notably, this technique is entirely training-free.
Specifically, as illustrated in Fig. 3, the initial step involves performing the standard denoising process for completing flow and caching flow attention. Subsequently, starting from step $t-1$, the noisy latent $\mathbf{z}$ is divided into two subsets based on parity. Latent interpolation process then follows a two-step alternating loop: even-indexed latents (shown in red) undergo denoising, while odd-indexed latents (shown in green) are obtained by warping operations using bi-directional optical flows. In the next step, only the interpolated latents (green) are denoised, and the red latents are generated through a similar warping process. Due to the negligible time cost of warping latent, the latency of denoising is halved by processing only half of the frame latent at each sampling timestep. Notably, the warping operation needs to be conducted at the $\mathbf{z}_{0}$ stage (as per Equation 4).
Figure 5: Qualitative comparisons. We compare our method against several approaches on BR and OR tasks.
Ideally, this process could be iterated until the final denoising step. However, since we are operating in latent space, completed optical flow provides only coarse-grained guidance, corresponding to the early stage of the diffusion process. Therefore, we restrict latent interpolation to the initial $S$ denoising steps, during which the overall structure of the image is established [13]. Additionally, to minimize flow errors, we perform warp operations exclusively between adjacent frames. Furthermore, to mitigate potential occlusion issues in flow warping [11], we perform the copy-paste operation in each denoising step (see Algorithm in appendix).
5 Experiments
5.1 Implementation details.
FloED is built upon Stable Diffusion (SD) [1] and incorporates motion modules initialized from AnimateDiff v3 [7]. We employ RAFT [24] for flow estimation and SD v2 inpainting [2] for the anchor frame strategy.
For training, we utilize the video dataset from Open-Sora-Plan [14], which comprises over 400K high-quality videos sourced from Pexels and Pixabay platforms. During the training phase, we use 16 frames at a resolution of 256×256, and synthetic random masks are applied throughout the training process.
For the two-stage training process, the number of epochs is set to 5 and 30, respectively. And $\lambda$ is set to 0.1 during the second stage. A batch size of 128 is achieved through gradient accumulation across eight NVIDIA A800 GPUs.
During inference, we empirically define the speed-up step $S$ as 5 (25 steps). We have developed an evaluation benchmark comprising 100 previously unseen videos, with 50 designated for object removal (OR) and 50 for background restoration (BR). For BR task, we use synthetic random masks that focus on the background. For the OR task, object masks are obtained by applying Segment-Anything [12] (SAM) to each frame.
5.2 Qualitative Results
To comprehensively demonstrate the capabilities of our method, we evaluated FloED on real videos for object removal and background restoration tasks. As illustrated in Fig. 4, FloED effectively handles both OR and BR tasks across a wide range of mask sizes, including those covering more than half of the frame area. Our approach adeptly in-paints the specified regions with coherent and contextually appropriate content. Additionally, FloED exhibits excellent temporal consistency throughout the video.
5.3 Comparisons
We present a comprehensive evaluation of our method compared to other open-source video inpainting solutions, including VideoComposer [26], CoCoCo [38], and Propainter [37]. Building upon the same image inpainting [2] results from anchor frames, we also assess the performance of the representative I2V model Dynamicrafter [29] as another comparison. We generate the corresponding video prompts for BR using VideoGPT [18], while manually creating appropriate background prompts for OR. Following the prior method [37], we applied a copy-paste approach to replicate the out-of-mask region across all methods except Dynamicrafter [29].
Qualitative comparisons.
Fig. 5 (left) compares the performance of background restoration, while Fig. 5 (right) illustrates object removal. Compared to the representative transformer-based method Propainter [37], our approach generates more vivid and detailed results in object removal (right), whereas Propainter encounters difficulties when synthesizing new content (left). Regarding text-guided diffusion methods, CoCoCo [38] and VideoComposer [26] tend to produce hallucinated content that is inconsistent with the surrounding environment. When focusing on the shadow regions within the mask area (right), our method demonstrates superior temporal consistency compared to Dynamicrafter [29].
Task BR OR
Metrics PSNR ↑ SSIM ↑ VFID ↓ ${E_{warp}}$ TC ↑ TA ↑
DC 17.50 0.4157 0.934 6.17 0.985 22.44
VC 22.81 0.8614 0.193 3.43 0.987 21.30
Propainter 27.02 0.9057 0.129 1.51 0.996 20.94
CoCoCo 23.08 0.8694 0.165 3.73 0.991 21.97
Ours 29.17 0.9441 0.118 2.83 0.994 22.49
Table 1: Quantitative comparisons using different metrics, including PSNR, SSIM, FVID, Optical flow warping error, Temporal consistency (TC), and Text alignment (TA).
Figure 6: We conduct reliable User Study with randomized order to assess their performance across different inpaintng outcomes.
Quantitative comparisons.
(a) Metric Evaluation. For background restoration, we employ PSNR [28], VFID [25], and SSIM [27] to quantify basic quality. Additionally, we assess temporal consistency using flow warping error [6, 13] in conjunction with Temporal Consistency (TC) [36]. TC is measured by the cosine similarity between consecutive frames in the CLIP-Image [8] feature space. For object removal, since ground truth data is unavailable for evaluating the aforementioned metrics, we utilize Text Alignment (TA) [8] as an evaluation metric, leveraging the CLIP score. For consistency, all metric evaluations were conducted at a resolution of 512×512. As shown in Tab. 1, our method outperforms other methods in PSNR, SSIM, VFID, and TA. Additionally, it surpasses other diffusion-based methods in flow warp error and TC, second only to Propainter by a small margin, due to the inherent generative characteristics of diffusion. Overall, FloED delivers excellent temporal consistency and high-quality outcomes compared to the state-of-the-art solutions. (b) User study. Since CLIP scores do not always align with human perception [36], we conducted a comprehensive user study. 12 annotators evaluated different inpainting results of 100 videos from 5 methods, assessing temporal consistency, content coherence, and overall quality. As illustrated in Fig. 6, our model was highly favored, achieving the highest scores in both BR (50.0%) and OR (40.33%), thereby demonstrating a consistent advantage over competing approaches.
5.4 Ablation Study
FA AF PSNR ↑ SSIM ↑ VFID ↓ ${E_{warp}}$ TC ↑
21.30 0.8435 0.246 4.15 0.989
25.34 0.9138 0.195 3.87 0.984
27.05 0.9255 0.170 3.03 0.989
28.71 0.9401 0.125 2.95 0.990
29.17 0.9441 0.118 2.83 0.994
Table 2: Ablation study. FA, AF stand for Flow Adapter and Anchor frame, respectively. All variants were trained using identical settings, except the last one with extended training (30 epochs).
Figure 7: Optical Flow Completion. Flows in first row are estimated by RAFT [24]. The completed flows provide motion guidance aligned with the mask surroundings.
Figure 8: Analysis of flow adapter. OR prompts are ”Green leaves of the bush sway in the wind.” and ”A huge rock with flowing water in the background.” respectively.
Effect of flow guidance.
We present the completed flow for the BR and OR tasks, as shown in Fig. 7. The corrupted flow is contextually inpainted with appropriate content that aligns with the surrounding environment while keeping the out-of-mask region intact, which serves as motion guidance for the primary inpainting branch. In Fig. 8, we further demonstrate the effects of flow guidance through two object removal examples. We compare two different variants of our model ($3^{th}$ and $5^{th}$ rows of Tab. 2) to illustrate the impact of flow guidance on inpainting. The inpainted leaves with the flow adapter are more coherent with the surrounding environment (left), especially in the regions at the edges of the object mask. Additionally, the repaired rocks exhibit greater temporal consistency (right) compared to outcomes without the flow adapter which demonstrates flow adapter ensures flow guidance effectively assists the model in generating content consistent with the environment, thereby enhancing temporal coherence and overall quality. Furthermore, as demonstrated in $3^{th}$ and $5^{th}$ rows of Tab. 2, the flow adapter significantly improves FloED’s performance across all metrics, providing substantial benefits to video quality and temporal consistency.
Effect of anchor frame.
As illustrated in Fig. 9, we investigate the impact of the anchor frame. Compared to the variant without the anchor frame strategy, incorporating an anchor frame introduces finer details, thereby enhancing the performance of FloED as an auxiliary strategy. Compared to the flow adapter, its effectiveness is somewhat inferior. As shown in $3^{th}$ and $4^{th}$ rows of Tab. 2, the variant incorporating the flow adapter achieves higher results than the one trained with anchor frame strategy.
Figure 9: Analysis of anchor frame. BR prompt is ”Grilled chicken wings on the barbecue, 8K. ”
Figure 10: Speeding steps study. Performance markedly deteriorates when the acceleration step surpasses five.
Flow Branch L.I. Flow Cache Time SSIM↑
15.55 0.843
$1^{st}$ step 19.79 0.945
$1^{st}$ step $2^{nd}\sim 6^{th}$ 17.72 (↓ 10.5%) 0.944
$1^{st}$ step $2^{nd}\sim 6^{th}$ $6^{nd}\sim 25^{th}$ 16.06 (↓ 18.8%) 0.944
Table 3: Efficiency analysis. L.I. represents the Latent Interpolation. The second row represents no flow-related computation involved. As the closest counterpart, CoCoCo [38] takes 26.25s.
Efficiency analysis.
We conducted an efficiency analysis using a video composed of 16 frames with an A800 GPU. The denoising process is executed for 25 steps with the classifier-free guidance scale set to 15. As presented in Tab. 3, we compare different efficiency strategies against the variant without any flow-related module. As previously mentioned, since our flow branch is independent of time-step during training, during the testing phase, we only need to utilize the flow branch in the first step of the denoising process to complete the damaged optical flow and cache the memory bank. For the remaining denoising steps, we can directly use the completed flow for latent interpolation and the cached K, V for flow guidance. As discussed in Sec. 4.4, we apply latent interpolation exclusively during the initial phase. Fig. 10 illustrates that continuously increasing the acceleration step beyond the early stage of the denoising process results in a sharp decline in performance. By utilizing flow-guided latent interpolation to accelerate the initial denoising and incorporating flow caching in subsequent phases, we can minimize denoising time with only a slight compromise in performance.
Thus, we determined an optimal solution: applying latent interpolation for the initial 5 steps ($2^{nd}\sim 6^{th}$) and flow caching for the remaining steps, resulting in 18.8% speed-up. Compared to the pure variant, which does not incorporate flow completion and flow attention, these efficiency benefits nearly offset the additional computational burden, incurring minimal cost. Notably, FloED outperforms its closest counterpart CoCoCo [38] by 10.19s in total.
Discussion.
Since no trainable modules are involved, latent interpolation can be directly adapted to diffusion-based methods like CoCoCo [38] for acceleration. We acknowledge that obtaining the completed optical flow beforehand may limit its transferability. Additionally, FloED primarily targets BR and OR tasks but is also capable of video editing. Meanwhile, compared to the conventional flow-guided method Propainter [37], FloED demonstrates superior performance in OR during flow completion. These points are further discussed in the appendix.
6 Conclusion
In this paper, we introduce FloED, an advanced video inpainting framework that efficiently incorporates flow guidance into diffusion models. By employing a dual-branch architecture, FloED first completes corrupted flow through a dedicated flow branch, followed by a multi-scale flow adapter that provides effective motion guidance to the primary inpainting branch. Furthermore, we developed a training-free latent interpolation technique that accelerates the multi-step denoising process. Complemented by a flow attention cache mechanism, FloED minimizes the additional computational burden introduced by incorporating optical flow. Comprehensive experiments demonstrated that FloED outperformed state-of-the-art methods in both background restoration and object removal tasks, showcasing its superior ability to maintain temporal consistency and content coherence in video inpainting.

### v3

@@
-Advanced Video Inpainting Using Optical Flow-Guided Efficient Diffusion
+Coherent Video Inpainting Using Optical Flow-Guided Efficient Diffusion
 Abstract
-Recently, diffusion-based methods have achieved great improvements in the video inpainting task.
-However, these methods still face many challenges, such as maintaining temporal consistency and the time-consuming issue.
-This paper proposes an advanced video inpainting framework using optical Flow-guided Efficient Diffusion, called FloED.
-Specifically, FloED employs a dual-branch architecture, where a flow branch first restores corrupted flow and a multi-scale flow adapter provides motion guidance to the main inpainting branch.
-Additionally, a training-free latent interpolation method is proposed to accelerate the multi-step denoising process using flow warping.
-Further introducing a flow attention cache mechanism, FLoED efficiently reduces the computational cost brought by incorporating optical flow.
-Comprehensive experiments in both background restoration and object removal tasks demonstrate that FloED outperforms state-of-the-art methods from the perspective of both performance and efficiency.
+The text-guided video inpainting technique has significantly improved the performance of content generation applications.
+A recent family for these improvements uses diffusion models, which have become essential for achieving high-quality video inpainting results, yet they still face performance bottlenecks in temporal consistency and computational efficiency.
+This motivates us to propose a new video inpainting framework using optical Flow-guided Efficient Diffusion (FloED) for higher video coherence.
+Specifically, FloED employs a dual-branch architecture, where the time-agnostic flow branch restores corrupted flow first, and the multi-scale flow adapters provide motion guidance to the main inpainting branch.
+Besides, a training-free latent interpolation method is proposed to accelerate the multi-step denoising process using flow warping.
+With the flow attention cache mechanism, FLoED efficiently reduces the computational cost of incorporating optical flow.
+Extensive experiments on background restoration and object removal tasks show that FloED outperforms state-of-the-art diffusion-based methods in both quality and efficiency.
+Our codes and models will be made publicly available.
 footnotetext: *Work during DAMO Academy internship. $\dagger$ Co-corresponding author.
 1 Introduction
-Video inpainting aims to predict corrupted regions by filling them with contextually appropriate and temporally coherent content, which plays an essential role in computer vision, particularly in background restoration (BR) [23, 17, 34] and object removal (OR) [4, 33, 19].
-Although conventional transformer-based approaches [15, 5, 37] have made great progress on standard benchmarks [31, 20], they encounter difficulties when requiring the synthesis of new content not present in existing frames.
-Currently, diffusion models [36, 38] have demonstrated significant prowess in text-guided video inpainting.
-However, temporal consistency remains an area with substantial scope for further improvement.
-And current methods struggle to generate satisfying results on BR and OR tasks, which is probably caused by the training priors of powerful base models.
-As demonstrated in Fig.
-1, contents inpainted by CoCoCo [38] lacks harmony with its surroundings.
-This disharmony can be alleviated by providing motion guidance that is aligned with the surroundings.
-Optical flow, a key modality for motion information, provides significant guidance and enhances temporal consistency.
-Notably, integrating optical flow into video inpainting requires additional operations, including flow completion and effective incorporation, which incur extra computational costs.
-And diffusion inherently suffers from efficiency due to the multi-step denoising process.
-Consequently, when leveraging optical flow, it is essential to consider efficiency enhancements tailored to the multi-step characteristics of diffusion models.
-Figure 1: Superiority of FloED on background restoration and object removal.
-Based on the above analysis, we propose an advanced video inpainting framework using optical Flow-guided Efficient Diffusion, called FloED.
-By leveraging motion information, our approach aims to enhance the reliability and performance of diffusion-based techniques in BR and OR applications.
-Initially, built on the large-scale T2I model, we inflate the origin layers and integrate pre-trained motion module [7] to establish a basic text-to-video framework.
-Then we fine-tune the motion module to effectively adapt it for the video inpainting domain.
-By adopting this framework as the primary inpainting branch, our innovations focus on three key aspects: (1) We design a tailored flow branch that completes corrupted flow while maintaining consistent channel numbers with the primary branch.
-Next, we integrate a multi-scale flow adapter that incorporates flow features into the decoder blocks of the primary U-Net architecture, enabling FloED to utilize motion information more effectively.
-(2) Building on the observation that adjacent latents share similar motion patterns [35] and recognizing that the essence of diffusion models involves fundamentally multi-step sampling processes, we introduce a training-free latent interpolation technique.
-This method leverages warping operation guided by optical flow to effectively accelerate the multi-step denoising process during early denoising stage.
-Furthermore, by incorporating a flow attention cache mechanism during the remain denoising stage as a complementary speed-up solution, we efficiently minimize the additional computational burden typically introduced by flow adapters and flow branches.
+Text-guided video inpainting aims to predict corrupted regions with text-aligned and temporally coherent contents, which has drawn widespread attention and applications.
+Conventional transformer-based methods [4, 17, 15, 31, 34] fundamentally lack textually semantic controllability when users require precise semantic alignment between textual descriptions and generated video results.
+In recent years, diffusion models [21, 9, 20, 19] have demonstrated extraordinary capability in generating realistic and text-aligned content.
+Thus, diffusion-based solutions [10, 28, 35] have made substantial progress in the realm of text-guided image inpainting.
+However, directly applying these methods to video inpainting falls short in maintaining the necessary temporal consistency.
+Although rapid developments in text-to-video (T2V) generative diffusion models [6, 27] have successfully facilitated text-guided video inpainting [33, 36, 14] by leveraging motion modules [6] for temporal consistency, it still remains an area with substantial scope for further improvement.
+Meanwhile, current diffusion-based solutions struggle to produce satisfactory results in both Background Restoration (BR) and Object Removal (OR) scenarios.
+Both BR and OR need to generate background content semantically aligned with textual descriptions, while ensuring spatial-temporal coherence between synthesized content and contextual video semantics.
+As evidenced in Fig.
+LABEL:fig:headfig, state-of-the-art methods like CoCoCo [36] exhibit spatial-temporal disharmony where inpainted regions demonstrate incompatible texture patterns or lighting conditions with surrounding contexts.
+We further hypothesize that this disharmony can be mitigated by integrating motion guidance that aligns with the scene.
+In this regard, optical flow, a key modality for capturing motion information, provides valuable guidance and enhances temporal consistency, potentially alleviating the observed inconsistencies.
+Notably, integrating optical flow into video inpainting requires additional operations, including flow estimation and flow completion and effective incorporation, all of which incur extra computational costs.
+Thus, diffusion models inherently suffer from efficiency due to the multi-step denoising process.
+Consequently, when we leverage optical flow, it is essential to consider efficiency enhancements tailored to the multi-step characteristics of diffusion models.
+Based on the above analysis, we propose a coherent video inpainting framework using optical Flow-guided Efficient Diffusion, called FloED.
+By leveraging motion information, our approach seeks to enhance both the performance and efficiency of diffusion-based techniques in BR and OR applications.
+Based on the Animatediff [6], we fine-tune the motion module in the first stage to effectively align its temporal modeling capacity with the video inpainting task.
+Built on this primary inpainting branch, our innovations focus on three key aspects: (1) We design a time-agnostic flow branch that completes corrupted flow while maintaining consistent channel numbers with the primary branch.
+Next, we integrate the multi-scale flow adapters that inject flow features into the decoder blocks of the primary U-Net architecture, enabling FloED to utilize motion information more effectively.
+(2) Building on the observation that adjacent latent features share similar motion patterns [31] and diffusion models fundamentally involve multi-step sampling processes, we introduce a training-free latent interpolation technique, which leverages warping operation guided by optical flow to effectively accelerate the multi-step denoising process during early denoising stage.
+Furthermore, by incorporating the flow attention cache mechanism during the remain denoising stage as a complementary speed-up solution, we efficiently minimize the additional computational burden typically introduced by flow adapters and the flow branch.
 (3) Recognizing that state-of-the-art image inpainting models significantly outperform video inpainting diffusion models, we utilize an anchor frame strategy to enhance the quality of video inpainting outcomes.
@@
 Our main contributions are as follows:
-  • We propose a dedicated dual-branch architecture that incorporates motion guidance with a multi-scale flow adapter, thereby enhancing temporal consistency and overall quality.
+  • Novel video inpainting model.
+We propose a dedicated dual-branch architecture that integrates optical flow guidance through flow adapters, thereby enhancing spatial-temporal consistency with compatible outcomes.
+  • Efficient denoising process.
 We introduce a training-free latent interpolation technique that leverages optical flow to speed up the multi-step denoising process.
-Complemented by a flow attention cache mechanism, FloED efficiently reduces the additional computational costs introduced by the flow.
-  • We conducted extensive experiments, including both quantitative and qualitative evaluations, to validate that FloED demonstrates superiority over other state-of-the-art methods in both BR and OR tasks.
-2 Related Work
-2.1 Conventional Video Inpainting
-Conventional video inpainting mainly focuses on two types of tasks: object removal and background restoration.
-Object removal [4, 33, 19] aims to eliminate unwanted objects from video frames.
-And background restoration [23, 17, 34] involves seamlessly reconstructing missing region of the background with coherent content.
-Given the fact that it’s more simple to complete flows instead of directly filling masked regions [32], conventional transformer-based models [15, 35, 37] primarily rely on optical flow to propagate features or pixels for better inpainting outcomes.
-Specifically, E2FGVI [15] introduces an end-to-end framework with flow-guided feature propagation.
-FGT [35] combines decoupled spatiotemporal attention with a flow-guided content propagation.
-ProPainter [37] advances this field by merging dual-domain propagation with a mask-guided transformer.
-Although propainter [37] has shown significant improvements on standard benchmarks [20, 31], it still struggles with generating coherent new content that isn’t present in existing frames.
-There is a pressing need to develop more robust solutions.
-2.2 Diffusion-based Video Inpainting
-In recent years, diffusion models [22, 9] have revolutionized the field of content generation, showcasing exceptional abilities in creating highly realistic outputs [21, 2].
-The development of text-to-video (T2V) [7, 10] and image-to-video (I2V) [30, 3] diffusion generative models, has paved the way for innovations in text-guided video inpainting.
-Models like AVID [36] and CoCoCo [38], have highlighted the routine practice of employing pre-trained Stable Diffusion Inpainting models [2] equipped with motion modules [7] for text-guided video inpainting.
-Specifically, AVID [36] introduce zero-shot generation pipeline with a structure guidance module.
-CoCoCo [38] implements enhanced attention mechanisms to address text-video alignment and ensure motion consistency across frames.
-While these advancements have contributed to progress, existing diffusion-based video inpainting methods still exhibit significant room for improving temporal consistency and they perform suboptimally on both OR and BR tasks.
-Compared with them, FloED showcases distinct advantages over temporal consistency and overall quality.
+Complemented by the flow attention cache mechanism, FloED efficiently reduces the additional computational costs introduced by the flow.
+  • State-of-the-art performance.
+We conducted extensive experiments on OR and BR tasks, including both quantitative and qualitative evaluations, to validate that FloED outperforms other state-of-the-art text-guided diffusion methods in terms of both performance and efficiency.
 Figure 1: Overview of FloED.
@@
 In the first training stage, we focus exclusively on the upper branch, optimizing the motion layer to adapt specifically to the video inpainting domain.
-Subsequently, we introduce a dedicated flow branch complemented by a multi-scale flow adapter, which provides flow guidance covering upblocks of primary UNet.
+Subsequently, we introduce a time-agnostic flow branch complemented by a multi-scale flow adapter, which provides flow guidance covering upblocks of primary UNet.
 During the inference phase, we enhance efficiency by integrating the flow attention cache (right part).
+2 Related Work
+Text-guided video inpainting, an emerging topic for synthesizing semantically consistent and temporally stable content in corrupted video regions through natural language instructions, has garnered significant research interest [33, 36, 14] and immense potential for creative and practical industrial applications.
+Although conventional transformer-based solutions [15, 31, 34] have shown significant improvements on standard benchmarks DAVIS [18] and Youtube-vos [29], they can’t provide semantic controllability when users require precise semantic alignment between textual descriptions and generated video results.
+In recent years, diffusion models [21, 9] have revolutionized the field of content generation, showcasing exceptional abilities in creating realistic outputs [20, 19] and controllable outcomes [30, 32].
+The application of diffusion models to text-guided image inpainting has seen remarkable advancements [10, 28, 35].
+However, directly applying these methods to video inpainting falls short in maintaining the necessary temporal consistency.
+The emergence of text-to-video technologies [6, 27, 3], has provided the solutions for the text-guided video inpainting.
+Notably, Animatediff [6] preserves temporal coherence in video generation through motion module training while effectively utilizing Stable Diffusion’s robust framework [1].
+Building upon this foundation, recent approaches [33, 36, 14] have advanced text-guided video inpainting.
+Specifically, AVID [33] proposes a fine-grained generation pipeline with structure guidance.
+CoCoCo [36] employs enhanced attention mechanisms to improve text-video alignment and motion consistency.
+And DiffuEraser [14] introduces the prior model for better inpainting outcomes.
+Nevertheless, existing methods demonstrate suboptimal performance on both object removal (OR) and background restoration (BR) tasks, frequently producing incompatible texture patterns or illumination conditions with surrounding contexts.
+In comparison, FloED exhibits superior temporal consistency and overall output quality.
 3 Preliminaries
@@
 where $\epsilon_{\theta}(\mathbf{z}_{t},t,c)$ is the predicted noise based on $\mathbf{z}_{t}$, the time step $t$ and the condition $c$.
-Once trained, we could leverage the deterministic sampling of DDIM [9] and PNDM [16] to denoise $\mathbf{z}_{t}$:
+Once trained, we could leverage the deterministic sampling of DDIM [9] to denoise $\mathbf{z}_{t}$:
 $\mathbf{z}_{t-1}$ $=\sqrt{\alpha_{t-1}}\underbrace{\hat{\mathbf{z}}_{t\rightarrow 0}}_{\text{predicted `$\mathbf{z}_{0}$'}}+$ (3)
@@
 4 Methods
-Building upon Stable Diffusion inpainting model [1], we integrate the motion module from Animatediff [7], adopting it as primary inpainting branch of our framework.
-As illustrated in Fig.
-2, the training process of FloED can be divided into two stages.
-Initially, we adapt primary branch to video inpainting by fine-tuning the motion modules (Sec.
-4.1).
-Our approach then incorporates a dedicated flow branch to complete corrupted flows estimated from masked frames, alongside a multi-scale flow adapter that provides motion guidance for the primary inpainting branch (Sec.
+Figure 2: Illustration of flow-guided latent interpolation (left) and warping operation (right) during the denoising process.
+Given a text prompt $P$ and original video sequence $\mathbf{V_{0}}=\{\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{N-1}\}\in\mathbb{R}^{N\times 3\times H\times W}$ and a binary mask sequence $\mathbf{m}=\{\mathbf{m}_{0},\mathbf{m}_{1},\ldots,\mathbf{m}_{N-1}\}\in\mathbb{R}^{N\times 1\times H\times W}$, corrupted frames $\mathbf{V_{m}}$ are obtained by applying the Hadamard product as follows: $\mathbf{V_{m}}=\mathbf{V_{0}}\odot\mathbf{m}$.
+We aim to generate a set of spatiotemporally consistent inpainted outcomes with text-aligned contents which demonstrate compatible texture patterns and lighting conditions with surrounding context.
+4.1 Network Overview
+An overview of our proposed model, FLoED, is depicted in Fig.
+1.
+For the architecture, FloED adopt pretrained Stable Diffusion Inpainting backbone [1] as the primary branch, while incorporating motion modules initialized from AnimateDiff v3 [6].
+The training process of FloED can be divided into two stages.
+In the first stage, we fine-tune the motion modules to align their temporal modeling capacity with video inpainting.
+In the second stage, FloED incorporates a dedicated flow branch to complete corrupted flows estimated from masked frames, alongside multi-scale flow adapters that inject hierarchical motion information into the primary inpainting branch (Sec.
 4.2).
@@
 4.4).
-4.1 Domain Adaptation
-Given an original video sequence $\mathbf{V_{0}}=\{\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{N-1}\}\in\mathbb{R}^{N\times 3\times H\times W}$ and a binary mask sequence $\mathbf{m}=\{\mathbf{m}_{0},\mathbf{m}_{1},\ldots,\mathbf{m}_{N-1}\}\in\mathbb{R}^{N\times 1\times H\times W}$, corrupted frames $\mathbf{V_{m}}$ are obtained by applying the Hadamard product as follows: $\mathbf{V_{m}}=\mathbf{V_{0}}\odot\mathbf{m}$.
-We aim to generate a set of spatiotemporally consistent inpainted outcomes with temporally coherent and contextually appropriate content in the corrupted area while ensuring out-of-mask areas are unchanged.
-Following previous solutions [36, 38], FloED is built on Stable Diffusion (SD) inpainting model [1], equipped with motion module from AnimateDiff-v3 [7] for better temporal consistency.
-Notably, during the first stage, only the primary branch is taken into consideration, and we only optimize the motion layer for adapting to the video inpainting.
-Figure 3: Illustration of flow-guided latent interpolation (left) and warping operation (right) during the denosing process.
+Notably, noised video latent $\mathbf{V_{t}}$, down-sampled binary mask $\mathbf{m}$, corrupted video latent $\mathbf{V_{m}}$ are concatenated along the channel axis as input.
 4.2 Flow Guidance in Video inpainting
-Typically, for video inpainting methods [15, 5, 37] that utilize optical flow, it is typically necessary to first complete the corrupted flow because only an estimated flow from the masked frames is initially available.
-Specifically, we first employ a pre-trained flow estimator [24] to perform corrupted flow estimation from $\mathbf{V_{m}}$.
-Subsequently, we introduce a flow sub-branch composed of the first layer from each block of the primary branch, which is independent of the timestep.
-As illustrated in Fig.
-2, the flow sub-branch completes the corrupted flow while maintaining consistent channel numbers with the latent representations of the corresponding frames in the primary branch.
-This approach enables more effective utilization of motion information by integrating multi-scale flow adapters into all upblocks of the primary UNet.
-As for the flow adapter, it consists of learnable modules with decoupled cross-attention mechanisms that embed flow features into the primary branch (Fig.
-2), thereby providing robust motion guidance effectively.
-We choose to insert the flow adapter between the text attention and the motion layer because it can compensate for the information provided by the prompt and further supply temporal guidance to the motion layer.
-Consequently, in the second training phase, we also trained the motion layer simultaneously.
-An additional optical loss $\mathcal{L}_{\textrm{flow}}$ is introduced in the second stage: ($F$ represents the ground truth flow.)
+Considering the flow guidance, FloED first employs a pre-trained flow generator RAFT [22] to obtain the corrupted flow from $\mathbf{V_{m}}$, which is annotated as ”Corrupted Flow Estimation” in Fig.
+1.
+Flow Completion Branch.
+To reconstruct the corrupted flow, the framework incorporates a dedicated flow completion branch that architecturally aligns with the primary inpainting backbone.
+Specifically, it selectively aggregate the initial ResNet module [7] from each corresponding block in the primary branch, to ensure channel-wise compatibility.
+And we exclude the time-step from ResNet to obtain the time-agnostic flow completion branch which creates stable flow feature regardless of diffusion progression.
+As shown in Fig 1, this flow completion branch enables comprehensive motion guidance by injecting compatible motion information into up-blocks of the primary UNet branch with multi-scale flow adapters.
+Flow Adapter.
+Our flow adapter is inspired by IP-Adapter [30], which is consisted with a separated cross-attention layer.
+The reconstructed flow features are fed into the cross attention for motion guidance.
+Notably, FloED strategically positions the flow adapter between the text cross-attention and motion modules to enable flow-aware latent modulation.
+This critical design further addresses misalignment caused by text-driven generation, because the multi-scale flow adapters dynamically adjusts latent features from the text cross-attention layer using optical flow priors, ensuring synthesized content maintains spatial-temporal compatibility with surrounding contexts.
+And in the second stage, we also continually fine-tune the motion modules.
+The reconstructed optical loss $\mathcal{L}_{\textrm{flow}}$ is introduced in the second stage. ($F$ represents the ground truth flow)
 $\mathcal{L}_{\textrm{sec}}=\mathcal{L}_{\textrm{diff}}+\lambda*\mathcal{L}_{\textrm{flow}},\mathcal{L}_{flow}=\|\hat{F}-F\|_{1}$ (5)
 4.3 Anchor Frame Strategy
-Recognizing that state-of-the-art image inpainting models significantly outperform their video inpainting counterparts, we introduce an anchor frame strategy to enhance the quality of video inpainting results.
+Recognizing that state-of-the-art image inpainting models developed in academia far surpass their video inpainting counterparts, we introduce an anchor frame strategy that leverages these advanced image techniques to substantially improve video inpainting performance.
 As illustrated in Fig.
-2, for a given video sequence $\mathbf{V_{0}}$, we select an additional frame from the beginning of the sequence to serve as an anchor frame.
+1, for a given video sequence $\mathbf{V_{0}}$, we select an additional frame from the beginning of the sequence to serve as an anchor frame.
 We then utilize a pre-trained text-to-image (T2I) inpainting model [2] to reconstruct its corrupted region in advance.
@@
 This strategy leverages the superior performance of image inpainting models to improve the overall quality of video inpainting.
-Figure 4: Qualitative Results. The orange dotted line indicates the input source.
 4.4 Efficient Inference for FloED
@@
 This approach is complemented by a flow attention cache mechanism during the inference phase.
-Optical Flow Attention Cache Unlike the primary branch, the flow sub-branch is independent of timestep.
-During the inference phase, we utilize the flow sub-branch exclusively for flow completion in the first step and then use these completed flows for all subsequent steps.
-Regarding the flow adapter, it introduces additional computations by calculating the flow attention at every denoising step and multiple resolutions.
+Flow Attention Cache.
+Unlike the primary branch, the flow branch is independent of timestep.
+During the inference phase, we utilize the flow branch exclusively for flow completion in the first step and then use these completed flows for all subsequent steps.
+Regarding the multi-scale flow adapters, they introduce additional computations by calculating the flow attention at every denoising step and multiple resolutions.
 To optimize this process, we establish a cache mechanism by computing the keys and values only during the first step and storing them in the memory bank (right part of Fig.
-2).
+1).
 For the remaining steps, the cached keys and values are directly retrieved from the memory bank, eliminating the need for repeated calculations and enhancing efficiency.
-Training-Free Denoising Speed-up Since adjacent feature latent exhibit similar motion patterns [35] and diffusion models generate high-level content early in the denoising process [13], we aim to speed-up the denoising process by interpolating latent using the completed flow.
+Training-Free Denoising Speed-up.
+Since adjacent feature latent exhibit similar motion patterns [31] and diffusion models generate high-level content early in the denoising process [12], we aim to speed-up the denoising process by interpolating latent using the completed flow.
 Notably, this technique is entirely training-free.
 Specifically, as illustrated in Fig.
-3, the initial step involves performing the standard denoising process for completing flow and caching flow attention.
+2, the initial step involves performing the standard denoising process for completing flow and caching flow attention.
 Subsequently, starting from step $t-1$, the noisy latent $\mathbf{z}$ is divided into two subsets based on parity.
-Latent interpolation process then follows a two-step alternating loop: even-indexed latents (shown in red) undergo denoising, while odd-indexed latents (shown in green) are obtained by warping operations using bi-directional optical flows.
+The latent interpolation process then follows a two-step alternating loop: even-indexed latents (shown in red) undergo denoising, while odd-indexed latents (shown in green) are obtained by warping operations using bi-directional optical flows.
 In the next step, only the interpolated latents (green) are denoised, and the red latents are generated through a similar warping process.
 Due to the negligible time cost of warping latent, the latency of denoising is halved by processing only half of the frame latent at each sampling timestep.
-Notably, the warping operation needs to be conducted at the $\mathbf{z}_{0}$ stage (as per Equation 4).
-Figure 5: Qualitative comparisons. We compare our method against several approaches on BR and OR tasks.
+Notably, the warping operation needs to be conducted at $\mathbf{z}_{0}$ (as per Equation 4).
+Figure 3: Qualitative comparisons. We compare FloED against diffusion-based SOTAs on BR and OR tasks.
 Ideally, this process could be iterated until the final denoising step.
@@
 Additionally, to minimize flow errors, we perform warp operations exclusively between adjacent frames.
-Furthermore, to mitigate potential occlusion issues in flow warping [11], we perform the copy-paste operation in each denoising step (see Algorithm in appendix).
+Furthermore, to mitigate potential occlusion issues in flow warping, we perform the copy-paste operation in each denoising step (see in appendix).
 5 Experiments
+BR OR
+PSNR ↑ SSIM ↑ VFID ↓ ${E_{warp}}$ TC ↑ TA ↑
+VC 22.81 0.8614 0.193 3.43 0.987 21.30
+CoCoCo 23.08 0.8694 0.165 3.73 0.991 21.97
+DiffuEraser 24.23 0.8583 0.218 2.98 0.984 19.19
+Ours 29.17 0.9441 0.118 2.83 0.994 22.49
+Table 1: Quantitative comparisons. Metrics including PSNR, SSIM, VFID, Optical flow warping error(${E_{warp}}$), Temporal consistency (TC), and Text alignment (TA).
 5.1 Implementation Details.
-FloED is built upon Stable Diffusion (SD) [1] and incorporates motion modules initialized from AnimateDiff v3 [7].
-We employ RAFT [24] for flow estimation and SD v2 inpainting [2] for the anchor frame strategy.
-For training, we utilize the video dataset from Open-Sora-Plan [14], which comprises over 400K high-quality videos sourced from Pexels and Pixabay platforms.
-During the training phase, we use 16 frames at a resolution of 256×256, and synthetic random masks are applied throughout the training process.
-For the two-stage training process, the number of epochs is set to 5 and 30, respectively.
-And $\lambda$ is set to 0.1 during the second stage.
-A batch size of 128 is achieved through gradient accumulation across eight NVIDIA A800 GPUs.
-During inference, we empirically define the speed-up step $S$ as 5 (25 steps).
-We have developed an evaluation benchmark comprising 100 previously unseen videos, with 50 designated for object removal (OR) and 50 for background restoration (BR).
+Dataset and Benchmark.
+We utilize the Open-Sora-Plan dataset [13], splitting videos at scene cuts to obtain 421,396 high-quality video clips paired with captions.
+We further developed an evaluation benchmark comprising 100 previously unseen videos sourced from Pexels and Pixabay platforms, with 50 designated for object removal (OR) and 50 for background restoration (BR).
 For BR task, we use synthetic random masks that focus on the background.
 For the OR task, object masks are obtained by applying Segment-Anything [11] (SAM) to each frame.
-5.2 Qualitative Results
-To comprehensively demonstrate the capabilities of our method, we evaluated FloED on real videos for object removal and background restoration tasks.
-As illustrated in Fig.
-4, FloED effectively handles both OR and BR tasks across a wide range of mask sizes, including those covering more than half of the frame area.
-Our approach adeptly in-paints the specified regions with coherent and contextually appropriate content.
-Additionally, FloED exhibits excellent temporal consistency throughout the video.
+Each video is manually selected to ensure a diverse range of motion amplitudes and camera movement speeds, while guaranteeing 4K resolution and 100 frames in total.
+For the caption, we generate the corresponding video prompts using VideoGPT [16], while manually revising appropriate background prompts for OR.
+Training and Inference Details.
+We employed a two stage training strategy with a resolution of 512 and 16-frame video sequences.
+And we generated mask sequences with random directions, and shapes to simulate BR and OR tasks.
+The first stage is trained on 8 NVIDIA A800 GPUs for 5 epochs with a batch size of 8.
+And second stage is trained on 8 NVIDIA A800 GPUs for 30 epochs with a batch size of 128, which is achieved through gradient accumulation the $\lambda$ is set to 0.1 during the second stage.
+During inference, we use DDIM [9] and empirically define the speed-up step $S$ as 5 (25 steps in total).
 5.2 Comparisons
-We present a comprehensive evaluation of our method compared to other open-source video inpainting solutions, including VideoComposer [26], CoCoCo [38], and Propainter [37].
-Building upon the same image inpainting [2] results from anchor frames, we also assess the performance of the representative I2V model Dynamicrafter [29] as another comparison.
-We generate the corresponding video prompts for BR using VideoGPT [18], while manually creating appropriate background prompts for OR.
-Following the prior method [37], we applied a copy-paste approach to replicate the out-of-mask region across all methods except Dynamicrafter [29].
+We present comprehensive comparisons with open-sourced text-guided diffusion-based approaches, including VideoComposer [24], CoCoCo [36] and DiffuEraser [14].
 Qualitative Comparisons.
-Fig.
-5 (left) compares the performance of background restoration, while Fig.
-5 (right) illustrates object removal.
-Compared to the representative transformer-based method Propainter [37], our approach generates more vivid and detailed results in object removal (right), whereas Propainter encounters difficulties when synthesizing new content (left).
-Regarding text-guided diffusion methods, CoCoCo [38] and VideoComposer [26] tend to produce hallucinated content that is inconsistent with the surrounding environment.
-When focusing on the shadow regions within the mask area (right), our method demonstrates superior temporal consistency compared to Dynamicrafter [29].
-Task BR OR
-Metrics PSNR ↑ SSIM ↑ VFID ↓ ${E_{warp}}$ TC ↑ TA ↑
-DC 17.50 0.4157 0.934 6.17 0.985 22.44
-VC 22.81 0.8614 0.193 3.43 0.987 21.30
-Propainter 27.02 0.9057 0.129 1.51 0.996 20.94
-CoCoCo 23.08 0.8694 0.165 3.73 0.991 21.97
-Ours 29.17 0.9441 0.118 2.83 0.994 22.49
-Table 1: Quantitative comparisons using different metrics, including PSNR, SSIM, FVID, Optical flow warping error, Temporal consistency (TC), and Text alignment (TA).
-Figure 6: We conduct reliable User Study with randomized order to assess their performance across different inpaintng outcomes.
-Quantitative comparisons.
-(a) Metric Evaluation.
-For background restoration, we employ PSNR [28], VFID [25], and SSIM [27] to quantify basic quality.
-Additionally, we assess temporal consistency using flow warping error [6, 13] in conjunction with Temporal Consistency (TC) [36].
-TC is measured by the cosine similarity between consecutive frames in the CLIP-Image [8] feature space.
-For object removal, since ground truth data is unavailable for evaluating the aforementioned metrics, we utilize Text Alignment (TA) [8] as an evaluation metric, leveraging the CLIP score.
-For consistency, all metric evaluations were conducted at a resolution of 512×512.
-As shown in Tab.
-1, our method outperforms other methods in PSNR, SSIM, VFID, and TA.
-Additionally, it surpasses other diffusion-based methods in flow warp error and TC, second only to Propainter by a small margin, due to the inherent generative characteristics of diffusion.
-Overall, FloED delivers excellent temporal consistency and high-quality outcomes compared to the state-of-the-art solutions.
-(b) User study.
-Since CLIP scores do not always align with human perception [36], we conducted a comprehensive user study.
-12 annotators evaluated different inpainting results of 100 videos from 5 methods, assessing temporal consistency, content coherence, and overall quality.
-As illustrated in Fig.
-6, our model was highly favored, achieving the highest scores in both BR (50.0%) and OR (40.33%), thereby demonstrating a consistent advantage over competing approaches.
-5.4 Ablation Study
+As demonstrated in Fig.
+3, VideoComposer, CoCoCo and DiffuEraser exhibit persistent limitations especially in OR tasks, frequently generating visual artifacts and content hallucinations that disrupt semantic consistency with scene context.
+In contrast, FloED inpaints the mask region with compatible contents and demonstrates precise text-conditioned generation capabilities, achieving superior temporal consistency and overall coherence in both BR and OR tasks.
+Figure 4: Optical flow related ablation studies. (F) ablation study demonstrates FloED conducts flow warping at noise $\epsilon$ instead of $\mathbf{z}_{0}$.
 FA AF PSNR ↑ SSIM ↑ VFID ↓ ${E_{warp}}$ TC ↑
@@
 Table 2: Ablation study. FA, AF stand for Flow Adapter and Anchor frame, respectively. All variants were trained using identical settings, except the last one with extended training (30 epochs).
-Figure 7: Optical Flow Completion. Flows in first row are estimated by RAFT [24]. The completed flows provide motion guidance aligned with the mask surroundings.
-Figure 8: Analysis of flow adapter. OR prompts are ”Green leaves of the bush sway in the wind.” and ”A huge rock with flowing water in the background.” respectively.
-Effect of flow guidance.
-We present the completed flow for the BR and OR tasks, as shown in Fig.
-7.
-The corrupted flow is contextually inpainted with appropriate content that aligns with the surrounding environment while keeping the out-of-mask region intact, which serves as motion guidance for the primary inpainting branch.
-In Fig.
-8, we further demonstrate the effects of flow guidance through two object removal examples.
-We compare two different variants of our model ($3^{th}$ and $5^{th}$ rows of Tab.
-2) to illustrate the impact of flow guidance on inpainting.
-The inpainted leaves with the flow adapter are more coherent with the surrounding environment (left), especially in the regions at the edges of the object mask.
-Additionally, the repaired rocks exhibit greater temporal consistency (right) compared to outcomes without the flow adapter which demonstrates flow adapter ensures flow guidance effectively assists the model in generating content consistent with the environment, thereby enhancing temporal coherence and overall quality.
-Furthermore, as demonstrated in $3^{th}$ and $5^{th}$ rows of Tab.
-2, the flow adapter significantly improves FloED’s performance across all metrics, providing substantial benefits to video quality and temporal consistency.
-Effect of anchor frame.
-As illustrated in Fig.
-9, we investigate the impact of the anchor frame.
-Compared to the variant without the anchor frame strategy, incorporating an anchor frame introduces finer details, thereby enhancing the performance of FloED as an auxiliary strategy.
-Compared to the flow adapter, its effectiveness is somewhat inferior.
-As shown in $3^{th}$ and $4^{th}$ rows of Tab.
-2, the variant incorporating the flow adapter achieves higher results than the one trained with anchor frame strategy.
-Figure 9: Analysis of anchor frame. BR prompt is ”Grilled chicken wings on the barbecue, 8K. ”
+Figure 5: We conduct reliable User Study with randomized order to assess inpainting outcomes of 4 methods.
+Quantitative Comparisons.
+We take metric evaluation and user studies to demonstrate quantitative comparisons.
+(1) Metric evaluation.
+For BR tasks, we employ the PSNR [26], VFID [23], and SSIM [25] to quantify basic quality.
+Additionally, we assess temporal consistency using flow warping error [5, 12] in conjunction with Temporal Consistency (TC) [33].
+TC is measured by the cosine similarity between consecutive frames in the CLIP-Image [8] feature space.
+For OM tasks, since ground truth data is unavailable for evaluating the aforementioned metrics, we utilize Text Alignment (TA) [8] as an evaluation metric, which also leverages the CLIP score.
+For consistency, all metric evaluations were conducted at a resolution of 512×512.
+As shown in Tab.
+1, FloED outperforms other methods in all the metrics, which demonstrates state-of-the-art performance.
+(2) User study.
+Since CLIP scores do not always align with human perception [33], we conducted a comprehensive user study in which 15 annotators evaluated the inpainting results across both BR and OR tasks (100 videos), assessing temporal consistency, text alignment, and context compatibility to select the best one among 4 methods.As illustrated in Fig.
+5, our model was highly favored, achieving the highest scores in both BR (62.27%) and OR (56.40%).
+5.3 Ablation Studies
+Flow-related Ablation Studies.
+We conduct flow-related ablation studies to validate the motion guidance, with experimental results shown in Fig. 4.
+(1) Flow completion.
+Using an object removal scenario as a case study, completed optical flow results demonstrate that the corrupted flow undergoes context-aware inpainting with spatially and temporally coherent content which maintains alignment with the surrounding environment (comparing B with C).
+And the reconstructed results further validate the completion capability of our time-agnostic flow branch.
+(2) Flow adapter.
+These reconstructed flows provide crucial motion guidance to the primary inpainting branch, effectively enhancing context compatibility and achieving higher video coherence through multi-scale flow adapters.
+Comparing the outcomes of D with E, black artifacts and incompatible lighting condition can be largely reduced.
+The experimental findings demonstrate that the multi-scale flow adapters’ injection of motion guidance significantly improves environmental consistency in generated content, which in turn enhances temporal coherence and overall quality.
+Meanwhile, we also conduct the quantitative architecture ablation study in Tab.
+2.
+Compared with anchor frame strategy, the multi-scale flow adapters demonstrate superior efficacy in enhancing framework’s performance, confirming their critical role in FloED.
+More ablation study results are provided in the Appendix.
+(3) Flow warping operation.
+Instructively, our ablation study F reveals that applying flow warping to the intermediate noise estimate $\epsilon$ during latent interpolation leads to cumulative error propagation with blurriness results compared to baseline D.
+To circumvent this, we reformulate the warping operation in the clean latent space $\mathbf{z}_{0}$(as per Equation 4).
 Figure 6: Speeding steps study. Performance markedly deteriorates when the acceleration step surpasses five.
-Flow Branch L.I. Flow Cache Time SSIM↑
-15.55 0.843
-$1^{st}$ step 19.79 0.945
-$1^{st}$ step $2^{nd}\sim 6^{th}$ 17.72 (↓ 10.5%) 0.944
-$1^{st}$ step $2^{nd}\sim 6^{th}$ $6^{nd}\sim 25^{th}$ 16.06 (↓ 18.8%) 0.944
-Table 3: Efficiency analysis. L.I. represents the Latent Interpolation. The second row represents no flow-related computation involved. As the closest counterpart, CoCoCo [38] takes 26.25s.
-Efficiency analysis.
-We conducted an efficiency analysis using a video composed of 16 frames with an A800 GPU.
-The denoising process is executed for 25 steps with the classifier-free guidance scale set to 15.
+Flow Branch Latent Interpolation Flow Cache Average time per frame(s)
+0.1287
+$1^{st}$ step 0.1491
+$1^{st}$ step $2^{nd}\sim 6^{th}$ 0.1342 (↓ 9.9%)
+$1^{st}$ step $2^{nd}\sim 6^{th}$ $6^{nd}\sim 25^{th}$ 0.1291 (↓ 13.4%)
+Table 3: Efficiency ablation study (432 $\times$ 240). The second row represents variant with no flow-related module involved (blue).
+Efficiency.
+In this section, we conduct the efficiency experiments with single H800 NVIDIA GPU under FP16 setting.
+And FloED’s denoising process is executed for 25 steps with the classifier-free guidance scale (CFG) $>$ 1.
+(1) Latent interpolation steps.
+As discussed in Sec.
+4.4, we apply latent interpolation exclusively during the initial phase.
+Thus, we conduct a speeding steps study in Fig.
+6.
+As depicted in Fig.
+6, continuously increasing the acceleration step results in a sharp decline in performance when the speeding steps $S$ beyond the early stage of the denoising process.
+Our experimental findings demonstrate that utilizing flow-guided latent interpolation for the initial 5 steps, we can minimize denoising time with only a slight compromise in performance.
+(2) Efficiency ablation study.
 As presented in Tab.
@@
 For the remaining denoising steps, we can directly use the completed flow for latent interpolation and the cached K, V for flow guidance.
-As discussed in Sec.
-4.4, we apply latent interpolation exclusively during the initial phase.
-Fig.
-10 illustrates that continuously increasing the acceleration step beyond the early stage of the denoising process results in a sharp decline in performance.
-By utilizing flow-guided latent interpolation to accelerate the initial denoising and incorporating flow caching in subsequent phases, we can minimize denoising time with only a slight compromise in performance.
-Thus, we determined an optimal solution: applying latent interpolation for the initial 5 steps ($2^{nd}\sim 6^{th}$) and flow caching for the remaining steps, resulting in 18.8% speed-up.
+Thus, we determined an optimal solution: applying latent interpolation for the initial 5 steps ($2^{nd}\sim 6^{th}$) and leveraging flow caching for the remaining steps for complement, resulting in 13.4% speed-up in the resolution of 432$\times$240.
 Compared to the pure variant, which does not incorporate flow completion and flow attention, these efficiency benefits nearly offset the additional computational burden, incurring minimal cost.
-Notably, FloED outperforms its closest counterpart CoCoCo [38] by 10.19s in total.
+(3) Efficiency comparisons.
+As shown the Tab.
+4, under the same denoising steps, FloED outperform other diffusion-based counterparts, such as CoCoCo and DiffuEraser, across all resolutions, demonstrating its state-of-the-art efficiency.
+Average time per frame (s) 512$\times$512 432$\times$256 256$\times$256
+Ours 0.3412 0.1291 0.0756
+CoCoCo 0.4033 0.2084 0.1735
+DiffuEraser 0.4176 0.5751 1.0336
+Table 4: Efficiency comparisons (CFG $>$ 1, 25 steps, FP16).
 Discussion.
-Since no trainable modules are involved, latent interpolation can be directly adapted to diffusion-based methods like CoCoCo [38] for acceleration.
-We acknowledge that obtaining the completed optical flow beforehand may limit its transferability.
-Additionally, FloED primarily targets BR and OR tasks but is also capable of video editing.
-Meanwhile, compared to the conventional flow-guided method Propainter [37], FloED demonstrates superior performance in OR during flow completion.
-These points are further discussed in the appendix.
+This paper focuses on text-guided video inpainting, primarily comparing with diffusion-based solutions.
+Our FloED also demonstrates superior performance over conventional flow-guided methods like ProPainter [34], with detailed comparisons provided in the Appendix.
+Furthermore, FloED’s latent interpolation can be directly extended to other diffusion-based approaches (e.g., CoCoCo [36]) for accelerated processing.
+However, we note that the strategy of pre-completing corrupted optical flow might restrict its transferability across different application scenarios.
 6 Conclusion
-In this paper, we introduce FloED, an advanced video inpainting framework that efficiently incorporates flow guidance into diffusion models.
-By employing a dual-branch architecture, FloED first completes corrupted flow through a dedicated flow branch, followed by a multi-scale flow adapter that provides effective motion guidance to the primary inpainting branch.
-Furthermore, we developed a training-free latent interpolation technique that accelerates the multi-step denoising process.
-Complemented by a flow attention cache mechanism, FloED minimizes the additional computational burden introduced by incorporating optical flow.
-Comprehensive experiments demonstrated that FloED outperformed state-of-the-art methods in both background restoration and object removal tasks, showcasing its superior ability to maintain temporal consistency and content coherence in video inpainting.
+In this paper, we introduced FloED, a coherent video inpainting framework that effectively integrates optical flow guidance into diffusion models to enhance temporal consistency and computational efficiency.
+By employing a dual-branch architecture, FloED first reconstructs corrupted flow, which then guides the inpainting process through multi-scale flow adapters.
+Additionally, our training-free latent interpolation technique and flow attention cache significantly reduce the computational overhead typically associated with optical flow integration.
+Experimental results demonstrate that FloED achieves state-of-the-art performance in both background restoration and object removal, showcasing its superior ability to maintain temporal consistency and content coherence in video inpainting.